A method for solving 5(x-2)-2(x-5)=9 is shown below.indentify the property used to obtain each of the 2 indicated steps. 5( x-2)-2(x-5)=9 #1 5x-10-2x+5=. #(2)5x-2x-10+10=9 3x+0=9. 3x=9x=3 how can I get toturin face to face on Saturday my phone. 516 2920918 thanks
Answer by ewatrrr(24785)

 
Hi,
Sry, we do not have the capability of face-to-face tutoring on this site
5( x-2)-2(x-5)=9
5x -10 -2x + 10=9 Distributive Property ('Distributing' the 5 and -2)
5x-2x-10+10=9 Cummutative Property ( 'Moving' Terms around )
3x = 9
x = 3
Migt recommend the following as a reference with examples:
http://www.purplemath.com/modules/numbprop.htm